# DiCeption: A Generalist Diffusion Model for Visual Perceptual Tasks

### v2

@@ B.1 Post-processing for Keypoints
B.2 Post-processing for RGB Masks
For entity segmentation and semantic segmentation RGB masks, we employ clustering algorithms to extract the object masks. Specifically, we first compute the histogram peaks for each of the three RGB channels and estimate the number of clusters by averaging the peak counts across the three channels. We then use KMeans clustering to group the colors and identify the clustered regions in the RGB mask. For each identified cluster, we extract regions with RGB values close to the cluster’s centroid. This step is followed by morphological operations to refine the extracted masks, such as filling holes and removing small, fragmented regions. We further filter the masks by computing their area, excluding any regions that are too small to be meaningful.
B.3 Performance Degradation of RGB Masks
@@
We observe that while the quality of our semantic segmentation visualizations is high, the average precision (AP) for certain categories remains unsatisfactory. For example, for the Person category, we conducted exhaustive experiments and achieved good visualization results (highlighted by the green rectangle in Figure S3), but AP is low (as in Table S2). Although, clearly, there is room for us to improve the semantic segmentation results, we do not intend to fit the data bias of those existing datasets, as pointed out by other authors, e.g., [87].
We trace the root cause of this issue to degradation during post-processing, particularly due to small objects and an excessive number of objects. Specifically, during mask processing, we filter out small noise regions, but this also removes some positive samples, such as the crowd and the bird highlighted in red in rows 3 to 5 in Figure S3. However, if we do not filter these noise regions, they further degrade the results. In our setting, filtering noise regions results in better metrics compared to not filtering them. Additionally, when an image contains an excessive number of objects of the same category (as in row 6 of Figure S3), post-processing may erroneously group similarly colored but distinct objects into a single class, leading to lower metrics. Furthermore, as in Table S2, we examine categories with fewer small objects and instances of those categories, such as bear, dog, and cat, and observe higher AP scores. However, for categories with opposite characteristics, their AP scores tend to be lower.
Although we can optimize post-processing for individual datasets by adjusting hyperparameters for each image to achieve the best results, this approach becomes impractical for large-scale in-the-wild evaluation, as it requires significant manual effort. Consequently, the dependency on post-processing remains a limitation of our approach.
@@
Appendix C Additional Results
C.1 Additional Visualizations
@@
We present additional visualization results of our method across various tasks, as can be seen in Figures S1S2S4S6S7S8S9S10S11S12S13S14. For point-prompted segmentation, we further compared our approach with SAM. These results strongly demonstrate the potential of DiCeption. DiCeption is capable of achieving high-quality results, even in challenging scenarios. Furthermore, the few-shot fine-tuning of DiCeption, which requires minimal data and trainable parameters, strongly demonstrates the remarkable transferability of DiCeptionto tackle new tasks. Our DiCeption is capable of further refining the segmentation of fine details, such as intricate hair structures.
@@
C.2 One-step Inference Does Not Work
Genpercept [114] demonstrates that one-step diffusion significantly enhances both the speed and accuracy of perceptual tasks. However, our experimental results reveal a notable increase in the proportion of failure cases when applying one-step diffusion in a multi-task setting, as illustrated in Figure S15. We believe that this is due to the potential overlap of denoising trajectories for different tasks. These overlapping trajectories can interfere with each other, resulting in failure cases. In contrast, multi-step denoising helps to mitigate such interactions. In a single-task setting, since the denoising trajectories pertain to a single task, this approach is more effective and stable.
Appendix D Limitations
Although our DiCeption achieves great results across multiple tasks, our model, as a diffusion model, leads to relatively longer inference times. On one H800, it takes an average of 0.8 seconds to process a single image. On one 4090-GPU card, inference for one image takes approximately 2 seconds. We believe that this issue can be addressed through few-step diffusion techniques. For tasks like semantic segmentation, the high demands of post-processing make the large-scale evaluation challenging. Despite the limitations, we believe that DiCeption is a valuable exploration for generalist visual foundation models.
@@ Appendix D Limitations
Figure S10: Additional point-prompted segmentation visualizations.
@@
Figure S14: Additional semantic segmentation visualizations.
Figure S15: The model tends to produce more failure cases in 1-step scenario.
Figure S16: A UNet-based model fails to perform multi-task.

### v3

@@ C.1 Post-processing for Keypoints
C.2 Post-processing for RGB Masks
@@
For entity segmentation and instance segmentation RGB masks, we employ clustering algorithms to extract the object masks. Specifically, we first compute the histogram peaks for each of the three RGB channels and estimate the number of clusters by averaging the peak counts across the three channels. We then use KMeans clustering to group the colors and identify the clustered regions in the RGB mask. For each identified cluster, we extract regions with RGB values close to the cluster’s centroid. This step is followed by morphological operations to refine the extracted masks, such as filling holes and removing small, fragmented regions. We further filter the masks by computing their area, excluding any regions that are too small to be meaningful.
C.3 Performance Degradation of Keypoints
[…]
C.4 Performance Degradation of RGB Masks
We observe that while the quality of our instance segmentation visualizations is high, the average precision (AP) for certain categories remains unsatisfactory. For example, for the Person category, we conducted exhaustive experiments and achieved good visualization results (highlighted by the green rectangle in Figure S9), but AP is low (as in Table S13).
We trace the root cause of metrics degradation during post-processing and find that this is particularly due to small objects and an excessive number of objects. Specifically, during mask processing, we filter out small noise regions. The genesis of these artifacts is predominantly attributed to subtle colorimetric fluctuations or minor inconsistencies in pixel values within areas of a mask intended to be uniformly colored. However, this operation also removes some positive samples, such as the crowd and the bird highlighted in red in rows 3 to 5 in Figure S9. These samples are susceptible to being misidentified as noise due to their diminutive size. Despite this limitation, the filtering of these noise regions is maintained because their persistence would otherwise exert a more detrimental impact on the quality of the final results. In our setting, filtering noise regions results in better metrics compared to not filtering them. Additionally, when an image contains an excessive number of objects of the same category (as in row 6 of Figure S9), post-processing may erroneously group similarly colored but distinct objects into a single class, leading to lower metrics. Furthermore, as in Table S13, we examine categories with fewer small objects and instances of those categories, such as bear, dog, and cat, and observe higher AP scores. However, for categories with opposite characteristics, their AP scores tend to be lower. This phenomenon is also observed in entity segmentation, which further elucidates why our entity segmentation results exhibit lower scores on small objects.
Although we can optimize post-processing by adjusting hyperparameters for each image to achieve the best results, this approach becomes impractical for large-scale evaluation, as it requires significant manual effort. Consequently, the dependency on post-processing remains a limitation of our approach.
@@ C.4 Performance Degradation of RGB Masks
Appendix D Additional Results
D.1 Additional Visualizations
@@
We present additional visualization results of our method across various tasks, as can be seen in Figures S8S11S10S15S16S17S18S19S20S21S22. For interactive segmentation, we compare our approach with SAM. These results strongly demonstrate the potential of DiCeption. DiCeption is capable of achieving high-quality results, even in challenging scenarios. Furthermore, the few-shot fine-tuning of DiCeption, which requires minimal data and trainable parameters, strongly demonstrates the remarkable transferability of DiCeption to tackle new tasks. Our DiCeption is capable of further refining the segmentation of fine details, such as intricate hair structures, achieving matting-level performance.
D.2 Comparative Experiments with One Diffusion
Qualitative visual comparisons between our method and One Diffusion in Figure S14 highlight key distinctions. In segmentation, our approach excels by simultaneously segmenting objects by semantic class and differentiating individual instances—a capability lacking in One Diffusion. Moreover, the segmentation quality of our method is superior to that of One Diffusion, especially in object-dense regions where the latter exhibits noticeable performance degradation.
A critical limitation of One Diffusion is its apparent inability to distinguish input images from conditioning signals, leading to a conflation of image understanding tasks with image generation. For example, when performing human keypoint estimation, One Diffusion may erroneously generate an image depicting a similar pose rather than predicting the actual keypoints. Conversely, our model, being fundamentally oriented towards image perception, not only consistently yields high-quality, accurate results without confusion, but also performs challenging perception tasks inaccessible to One Diffusion, such as interactive segmentation.
Appendix E Discussions and Limitations
Discussions
Our method highlights the following key findings:
  • The inherent prior knowledge of diffusion models is highly effective for perception tasks. By leveraging this prior effectively, our approach enables a single model to address multiple tasks. Notably, it achieves performance comparable to existing single-task specialized models, even on challenging tasks such as interactive segmentation, and does so with limited data.
  • Our comprehensive experimental evaluation demonstrates that token-wise concatenation is the most efficient and effective strategy for leveraging the prior knowledge of transformer-based diffusion models. Furthermore, we provide evidence that the DiT architecture works better compared to U-Net. This is attributed to the fact that transferring U-Net to multiple perception tasks not only introduces additional parameters that can potentially disrupt the pre-trained model’s prior but also suffers from significant information loss due to its inherent downsampling operations.
  • A modest CFG value can yield performance improvements for pixel-sensitive tasks such as depth and normal estimation.
  • We find that flow-matching models, when trained in a multi-step denoising setting, naturally support few-step inference for perception tasks.
  • Our DiCeption exhibits a faster and more effective adaptation capability to new downstream tasks.
  • The efficacy of our approach is demonstrated on a different DiT architecture and smaller model, indicating its robustness.
  • The model demonstrates strong capability of detail preservation after pixel-aligned training on perception tasks.
To the best of our knowledge, we are the first to successfully leverage diffusion priors to address multiple perception tasks with a single model without exceptionally large or cherry-picking high-quality data, achieving performance on par with specialized models, even on the challenging interactive segmentation compared with SAM. In our view, the capabilities of our method are far from being fully realized, and further training with larger, higher-quality datasets has the potential to yield even more compelling results. For instance, in high-level tasks such as referring segmentation shown in Figure S12, our model achieves results with finer details than the ground truth. This not only demonstrates the model’s ability to benefit from related tasks but also showcases its strong semantic understanding. Furthermore, we observe early signs of task composition in our model, albeit with a low success rate. For instance, the model can predict the depth or normal map of an object indicated by point inputs while generating a black mask for other regions, as illustrated in Figure S13, though the success rate is very low. In conclusion, we believe that our work not only presents a generalist model with a vast capacity for improvement, but also provides comprehensive experiments and analyses that can serve as a valuable foundation for future research.
Figure S12: DiCeption achieves finer results on referring segmentation, showing the potential of mutual improvement between related tasks.
Figure S13: Example of task composition. Our model can isolate a point-specified object to generate its corresponding depth map, while correctly suppressing predictions for all other regions. Although the success rate is very low, this result still reveals a promising capability.
Limitations
Although our DiCeption achieves great results across multiple tasks, our model, as a diffusion model, leads to relatively longer inference times. On one H800, it takes an average of 0.8 seconds to process a single image. On one 4090-GPU card, inference for one image takes approximately 2 seconds. We believe that this issue can be addressed through few-step diffusion techniques, which we leave for future works.
Furthermore, our evaluation on certain tasks such as human keypoints estimation and text-based instance segmentation necessitates post-processing, which can introduce substantial errors. However, unlike some contemporary diffusion-based works le2024diffusiongenerate ; wang2024lavin that often omit quantitative evaluation on the task such as human keypoints estimation, we take a step further by providing evaluation metrics. Our analysis demonstrates that lower scores on these tasks are not due to model performance but are significantly influenced by the post-processing step. Consequently, the dependence on post-processing for quantitative evaluation on certain tasks remains a limitation of our method. Despite the limitations, we believe that DiCeption is a valuable exploration for diffusion-based generalist visual perception foundation models.
@@ Appendix E Discussions and Limitations
Figure S18: Additional interactive segmentation visualizations.
@@
Figure S22: Additional text-based instance segmentation visualizations.